# RouteLLM: Learning to Route LLMs with Preference Data

### v3

@@ Abstract
Large language models (LLMs) exhibit impressive capabilities across a wide range of tasks, yet the choice of which model to use often involves a trade-off between performance and cost. More powerful models, though effective, come with higher expenses, while less capable models are more cost-effective. To address this dilemma, we propose several efficient router models that dynamically select between a stronger and a weaker LLM during inference, aiming to optimize the balance between cost and response quality. We develop a training framework for these routers leveraging human preference data and data augmentation techniques to enhance performance. Our evaluation on widely-recognized benchmarks shows that our approach significantly reduces costs—by over 2 times in certain cases—without compromising the quality of responses. Interestingly, our router models also demonstrate significant transfer learning capabilities, maintaining their performance even when the strong and weak models are changed at test time. This highlights the potential of these routers to provide a cost-effective yet high-performance solution for deploying LLMs.
$\dagger$\$$\dagger$\$footnotetext: Equal contribution.
@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language tasks. From open-ended conversation and question answering to text summarization and code generation, LLMs have showcased an impressive level of fluency and understanding [1, 9]. This rapid progress has been enabled by a combination of architectural innovations, such as the Transformer architecture [30], as well as improvements in scaling up data and training infrastructure [8, 26].
However, not all LLMs are created equal—there exists wide variation in the costs and sizes of LLMs, which can range in size from one billion to hundreds of billions of parameters. LLMs also differ in terms of the data they are trained on, which in turn leads to variations in the strengths, weaknesses, and capabilities of different models. Broadly speaking, larger models tend to be more capable but come at a higher cost, while smaller models tend to be less capable but cheaper to serve.
This heterogeneous landscape presents a dilemma in the practical deployment of LLMs in real-world applications. While routing all user queries to the largest, most capable model ensures high-quality results, it is prohibitively expensive. Conversely, routing queries to smaller models can save costs—up to more than 50x (e.g., Llama-3-70b vs. GPT-4, or Claude-3 Haiku vs. Opus [^1] Per one million output tokens: Llama-3-70b (\$1) vs. GPT-4-0613 (\$60), Haiku (\$1.25) vs. Opus (\$75))—but may result in lower quality responses, as the smaller model may not be able to handle complex queries effectively.
To this end, LLM routing is a promising solution to this problem, whereby each user query is first processed by a router model, before deciding which LLM to route the query to. This can potentially route easier queries to smaller models, and more difficult queries to larger models, optimizing the quality of model responses while minimizing cost. However, optimal LLM routing—defined as achieving the highest quality given a cost target or minimizing cost given a quality target—is a challenging problem. A robust router model needs to infer the intent, complexity, and domain of an incoming query as well as understand candidate models’ capabilities to route the query to the most appropriate model. Furthermore, the router model needs to be economical, fast, and adaptive to the evolving model landscape, where new models with improved capabilities are continually introduced.
@@
In this work, we present a principled framework for query routing between LLMs. Our setup involves learning to route between a stronger model and a weaker model as seen in Figure 1. Our objective is to minimize costs while achieving a specific performance target, such as 90% of the stronger model’s performance, by intelligently routing simpler queries to the weaker model and reserving more complex queries for the stronger model. We develop a training framework for router systems utilizing human preference data and data augmentation techniques. We evaluate our router models on widely recognized benchmarks, such as MMLU [17] and MT Bench [33], and demonstrate that our framework can significantly reduce costs—by over 2 times—without substantially compromising response quality.
@@
  • We formulate the LLM routing problem to explore the trade-off between cost and response quality.
  • We propose a router training framework based on human preference data and augmentation techniques, demonstrating over 2x cost saving on widely used benchmarks.
  • We open-source the code and preference data used to train our routers. [^2] https://github.com/lm-sys/RouteLLM.
Several recent studies have explored optimizing the cost and performance trade-offs in deploying large language models (LLMs). LLM-BLENDER [20] employs an ensemble framework that calls multiple LLMs at inference selecting the best response. Frugal-GPT [11] employs an LLM cascade, sequentially querying LLMs until a reliable response is found. AutoMix [2] relies on a smaller model self-verifying its response before potentially routing to a larger model. While these approaches rely on multiple LLM queries [18], our approach routes each query to a single LLM. FrugalML [10] focuses on prediction services with known labels rather than the general capabilities of LLMs. A closely related study, Hybrid-LLM [15], shares similarities with our framework but differs in three key ways: it uses synthetic preference labels derived via BARTScore [32], relies on a single BERT-based router architecture, and limits evaluation to in-domain generalization. In contrast, our work leverages human preference labels from Chatbot Arena [12], explores several router architectures, and demonstrates that augmenting the dataset results in significant performance improvements across all router architectures. Additionally, we emphasize out-of-domain generalization by evaluating on multiple public benchmarks.
@@ 2.1 Problem Formulation
Consider a set of $N$ different LLM models =$\mathcal{M}=\{M_{1},\dots,M_{N}\}$. Each model $M_{i}:\mathcal{Q}\rightarrow\mathcal{A}$ can be abstracted as a function that maps a query to an answer. A routing function $R:\mathcal{Q}\times\mathcal{M}^{N}\rightarrow\{1,\dots,N\}$ is an $N$-way classifier that takes a query $q\in\mathcal{Q}$ and selects a model to answer $q$, with the answer being $a=M_{R(q)}(q)$. The challenge of routing involves achieving an optimal equilibrium between increasing response quality and reducing cost [^3] We focus on inference cost in \$ cost/token, but the same concept applies to other costs such as latency.. Assume we have access to preference data: $\mathcal{D}_{\text{pref}}=\{(q,l_{i,j})\mid q\in\mathcal{Q},i,j\in N,l_{i,j}\in\mathcal{L}\}$, where $q$ is a query, and $l_{i,j}$ is a label representing the comparison outcome of comparing $M_{i},M_{j}$’s quality on $q$, which takes values in =$\mathcal{L}=\{\text{win}_{M_{i}},\text{tie},\text{win}_{M_{j}}\}$. It is important to distinguish between reward modeling [25] and routing. Reward modeling evaluates response quality post-LLM generation, whereas routing requires the router to select the appropriate model before seeing the response. This necessitates a deep understanding of the question’s complexity as well as the strengths and weaknesses of the available LLMs.
In this work, we focus on routing between two classes of models: (1) strong models ($\mathcal{M}_{\text{strong}}$), consisting of models capable of producing high quality responses but at a high cost. This class primarily consists of the most advanced closed-source models, such as GPT-4. (2) weak models ($\mathcal{M}_{\text{weak}}$), consisting of models with a relatively lower quality and lower cost, such as such as Mixtral-8x7B. This binary routing problem is quite common in practice, particularly as developers of LLM applications strive to balance quality and cost. Furthermore, addressing this problem forms the foundation for solving a more general $N$-way routing problem.
@@
We present a principled framework for learning a binary routing function $R_{\text{bin}}^{\alpha}:\mathcal{Q}\rightarrow\{0,1\}$ between $\mathcal{M}_{\text{weak}}$ and $\mathcal{M}_{\text{strong}}$ from preference data. To achieve this, we define $R_{\text{bin}}^{\alpha}$ using two components:
1) Win Prediction Model which predicts the probability of winning for strong models $\mathcal{M}_{\text{strong}}$, i.e. $P_{\boldsymbol{\theta}}(\text{win}_{\mathcal{M}_{\text{strong}}}|q)$. In our binary classification setting, this probability captures the win/loss probability of both model classes. [^4] We consider “tie” as a win for the weaker models. We can learn the parameters of this model $\boldsymbol{\theta}$ with maximum likelihood on the preference data:
$\max_{\boldsymbol{\theta}}\sum_{(q,l_{i,j})\in\mathcal{D}_{\text{pref}}}\log P_{\boldsymbol{\theta}}(l_{i,j}\mid q).$ (1)
By learning the winning probability on preference data, we capture the strengths and weaknesses of both model classes on various kinds of queries. In Section 3.2, we propose several approaches for parameterizing the win prediction model.
@@
2) Cost Threshold $\alpha\in[0,1]$ which converts the winning probability into a routing decision between $\mathcal{M}_{\text{strong}}$ and $\mathcal{M}_{\text{weak}}$. Given a query $q$, the routing decision is formulated as:
$R_{\text{bin}}^{\alpha}(q)=\begin{cases}0\text{ (i.e., $\mathcal{M}_{\text{weak}}$)}&\text{if }P(\text{win}_{M_{j}}\mid q)<\alpha,\\ 1\text{ (i.e., $\mathcal{M}_{\text{strong}}$)}&\text{otherwise}.\end{cases}$ (2)
The threshold $\alpha$ controls the quality/cost trade-off: a higher threshold imposes a stricter cost constraint, reducing expenses but potentially compromising quality.
Finally, we denote the router’s response as $\mathcal{M}_{R_{\text{bin}}^{\alpha}(q)}(q)$, which represents the response generated by either the weak or strong model, depending on the router’s decision.
@@ 2.2 Metrics
In this section, we define evaluation metrics that capture the trade-off between cost and quality in the LLM routing problem. We start with metrics that independently assess the quality and cost efficiency of a given $R_{\text{bin}}^{\alpha}$, then introduce two compounded metrics which we use in our experimental evaluations.
@@
For cost efficiency, we calculate the percentage of calls to the strong model:
$c(R_{\text{bin}}^{\alpha})$ $=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathbb{I}\{R_{\text{bin}}^{\alpha}(q)=1\},$ (3)
@@
since $\mathcal{M}_{\text{strong}}$ models incur significantly higher costs than $\mathcal{M}_{\text{weak}}$ models.
For quality, we measure the average response quality on an evaluation set $\mathcal{Q}$:
$r(R_{\text{bin}}^{\alpha})$ $=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\delta(\mathcal{M}_{R_{\text{bin}}^{\alpha}(q)}(q)),$ (4)
@@
where $\delta(\mathcal{M}_{R_{\text{bin}}^{\alpha}(q)}(q))$ represents a numerical score of the router’s response to $q$. This score can be result of a predefined metric measuring the correctness of the response in golden-labeled datasets (e.g., MMLU) or a numerical label (e.g., 1-5 or 1-10) where higher values indicate better quality.
@@
Given that the performance of $R_{\text{bin}}^{\alpha}$ lies between the weak and strong models’ performance, we quantify the router’s performance relative to the performance gap between both models. We define the overall performance gain of $R_{\text{bin}}^{\alpha}$ with the performance gap recovered (PGR):
$PGR(R_{\text{bin}}^{\alpha})=\frac{r(R_{\text{bin}}^{\alpha})-r(\mathcal{M}_{\text{weak}})}{r(\mathcal{M}_{\text{strong}})-r(\mathcal{M}_{\text{weak}})}.$ (5)
@@
Neither of these metrics alone is sufficient, as they do not capture the quality-cost trade-off in routing. For instance, a trivial router that sends every query to the strong model achieves a perfect $PGR=1$ without any cost reduction. Therefore, we compute a call-performance graph for a router $R_{\text{bin}}$ by varying the threshold values $\alpha$. We define the average performance gap recovered (APGR) as an overall measure of how well the router can recover the performance gap under different cost constraints:
$APGR(R_{\text{bin}})=\int_{0}^{1}PGR(R_{\text{bin}}^{\alpha})\;d\left(c(R_{\text{bin}}^{\alpha})\right).$ (6)
@@
In Figure 1-(right), APGR is represented by the area between the router’s performance curve and the weak model’s performance. Empirically, we discretize the percentage of calls interval $[0\%,100\%]$ into $\{c_{i}\}_{i\in[10]}$. For each $c_{i}$, we determine the cutoff threshold $\alpha_{i}$ that meets the cost constraint. We approximate $APGR$ using the following formula:
$APGR(R_{\text{bin}})\approx\frac{1}{10}\sum_{i=1}^{10}PGR(R_{\text{bin}}^{\alpha_{i}})$ (7)
@@
In many real-world applications, it is important to quantify the cost required to achieve a certain level of performance. Therefore, we define a second metric which we call call-performance threshold (CPT). Given a desired router performance (measured as PGR of $x\%$), the CPT($x\%$) refers to the minimum percentage of calls to the strong model required to obtained the desired PGR. In Figure 1-(right), the dotted green line denotes CPT(50%), i.e. the percentage of calls to GPT-4 required to achieve the desired performance of 50% PGR; here, $CPT(50\%)\approx 37\%$.

### v4

@@ Abstract
Large language models (LLMs) excel at a wide range of tasks, but choosing the right model often involves balancing performance and cost. Powerful models offer better results but are expensive, while smaller models are more cost-effective but less capable. To address this trade-off, we introduce a training framework for learning efficient router models that dynamically select between a stronger and weaker LLM during inference. Our framework leverages human preference data and employs data augmentation techniques to enhance performance. Evaluations on public benchmarks show that our approach can reduce costs by over 2 times without sacrificing response quality. Moreover, our routers exhibit strong generalization capabilities, maintaining performance even when routing between LLMs not included in training. This highlights the potential of our framework to deliver cost-effective, high-performance LLM solutions.
@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language tasks. From open-ended conversation and question answering to text summarization and code generation, LLMs have demonstrated an impressive level of fluency and understanding (Achiam et al., 2023; Bubeck et al., 2023). This rapid progress has been enabled by a combination of architectural innovations, such as the Transformer architecture (Vaswani et al., 2017), as well as scaling up data and training infrastructure (Brown et al., 2020; Radford et al., 2019).
However, not all LLMs are created equal—there exists wide variation in the sizes of different LLMs, which in turn affects the resources required to serve them. LLMs also differ in terms of the data on which they are trained, which in turn leads to variations in the strengths, weaknesses, and capabilities of different models. Broadly speaking, larger models tend to be more capable but come at a higher cost, while smaller models tend to be less capable but cheaper to serve.
This heterogeneous landscape presents a dilemma in the practical deployment of LLMs. Although routing all user queries to the largest and most capable model ensures high-quality results, it is prohibitively expensive. Conversely, routing queries to smaller models can save costs—by more than 50x (e.g., Claude-3 Haiku vs. Opus [^1] Per one million output tokens: Haiku (\$1.25) vs. Opus (\$75))—but may result in lower quality responses, as the smaller model may not handle complex queries effectively.
LLM routing (Ding et al., 2024; Hu et al., 2024) offers an effective solution by first processing each user query through a router, which then determines the most suitable LLM to handle the query. The router can direct simpler queries to smaller models and more complex ones to larger models, thereby balancing response quality with cost efficiency.
Achieving optimal LLM routing—maximizing quality within a cost constraint or minimizing cost for a target quality—is challenging. An ideal LLM router must (1) optimize response quality while invoking a single LLM per query, minimizing cost and latency as compared to multi-LLM approaches; (2) generalize to out-of-domain queries without needing separate routers for different domains; and (3) work across a broad range of LLMs without retraining, ensuring flexibility as the LLM landscape evolves.
@@
In this work, we introduce a principled framework for learning LLM routers from preference data. Our approach involves routing between two classes of models: (1) strong models, which provide high-quality responses at a high cost (e.g., GPT-4), and (2) weak models, which offer lower-quality responses at a reduced cost (e.g., Mixtral-8x7B). The objective is to minimize costs while achieving a specific performance target, e.g., 90% of the strong model, by intelligently routing simpler queries to a weak model and reserving more complex queries for the strong model. We use our framework to train several router models and evaluate them on widely recognized benchmarks such as MMLU (Hendrycks et al., 2020) and MT Bench (Zheng et al., 2023). We demonstrate that our router models significantly reduce costs—by over 2x—without substantially compromising quality. Moreover, they show strong performance across multiple strong / weak model pairs without requiring retraining.
@@
  • We propose a learning framework for routers that leverages human preference data and data augmentation techniques, achieving over 2x cost savings on popular benchmarks with minimal impact on response quality.
  • We demonstrate that our approach enables routers to generalize to unseen data while maintaining strong performance across multiple LLMs, allowing a single trained router to be effective across a wide range of use cases.
  • We open source our framework for training, serving, and evaluating LLM routers, allowing users to easily train their own routers and compare router performance across benchmarks.
2 Related Work
A key distinction exists between reward modeling (Ouyang et al., 2022) and LLM routing. Reward modeling assesses response quality after an LLM generates it, whereas routing involves selecting the appropriate LLM beforehand. This requires a deep understanding of the query’s complexity and the specific capabilities of available models.
Several recent works have also examined the cost-performance trade-offs in routing between different LLMs. LLM-Blender (Jiang et al., 2023) uses an ensemble framework that queries multiple LLMs during inference and selects the best response. Frugal-GPT (Chen et al., 2023) follows a cascading approach, sequentially querying LLMs until a reliable response is obtained. AutoMix (Aggarwal et al., 2024) uses a smaller model to self-verify its response before potentially routing the query to a larger model. These methods rely on multiple LLM queries, which can increase latency. In contrast, our approach routes each query to a single LLM, addressing the latency constraints of an ideal LLM router.
Hybrid-LLM (Ding et al., 2024) shares some similarities with our framework but differs in key aspects: it uses synthetic preference labels from the MixInstruct dataset (Jiang et al., 2023) based on BARTScore (Yuan et al., 2021) and relies on a single BERT-based router. In contrast, we leverage human preference labels from Chatbot Arena (Chiang et al., 2024) and explore multiple router architectures, showing that data augmentation significantly boosts performance across all architectures. Additionally, Hybrid-LLM evaluates on the MixInstruct test split and lacks evidence of out-of-domain generalization, whereas we aim to demonstrate this by evaluating on several decontaminated public benchmarks.
Finally, Zooter (Lu et al., 2023) uses routing labels from QwenRM reward models (Bai et al., 2023), which can inherit biases from their training data, affecting the reliability of the routing decisions. In contrast, our approach relies mainly on human preference data. Like Hybrid-LLM, Zooter explores only a BERT-style router. Additionally, their training signal relies on a fixed set of LLMs, limiting its adaptability to other LLMs. In contrast, we show that our approach maintains strong performance even with LLMs not included in the training data.
@@ 3.1 Problem Formulation
Consider a set of LLM models $\mathcal{M}$, where each model $M:\mathcal{Q}\rightarrow\mathcal{A}$ can be viewed as a function that maps a query $q\in\mathcal{Q}$ to an answer $a=M(q)\in\mathcal{A}$. In this work, we focus on routing between two classes of models: (1) strong models ($\mathcal{M}_{\text{strong}}$), which are capable of producing high-quality responses but come at a high cost, such as advanced proprietary models like GPT-4 (OpenAI, 2023), and (2) weak models ($\mathcal{M}_{\text{weak}}$), which offer lower-quality responses but at a reduced cost, such as models like Mixtral-8x7B (Jiang et al., 2024). This binary routing problem is common in practice, especially as users seek to optimize the trade-off between quality and cost by transitioning from closed-source to open-source models. Additionally, solving the binary routing challenge provides a foundation for extending to a more complex $N$-way routing scenario.
@@
Assume we have access to preference data: $\mathcal{D}_{\text{pref}}=\{(q,l_{s,w})\mid q\in\mathcal{Q},l_{s,w}\in L\}$, where $l_{s,w}$ represents the outcome of comparing the responses from a strong model $M_{s}\in\mathcal{M}_{\text{strong}}$ and a weak model $M_{w}\in\mathcal{M}_{\text{weak}}$ for a given query $q$, and takes values from the set $L=\{\text{win}_{s},\text{tie},\text{win}_{w}\}$. We introduce a principled framework for learning a binary routing function $R^{\alpha}:\mathcal{Q}\rightarrow\{\mathcal{M}_{\text{weak}},\mathcal{M}_{\text{strong}}\}$ from preference data. Our approach defines $R^{\alpha}$ using two key components:
1) Win Prediction Model: This model estimates the probability that a strong model in $\mathcal{M}_{\text{strong}}$ will outperform a weak model in $\mathcal{M}_{\text{weak}}$ for a given query $q$. This probability is denoted by $P_{\bm{\theta}}(\text{win}_{s}|q)$, where $\bm{\theta}$ represents the model parameters. These parameters are learned by maximizing the likelihood of the observed preference data:
$\max_{\bm{\theta}}\sum_{(q,l_{s,w})\in\mathcal{D}_{\text{pref}}}\log P_{\bm{\theta}}(l_{s,w}\mid q).$ (1)
By optimizing this likelihood, the model captures the comparative strengths and weaknesses of the two model classes across different query types. In Section 4.2, we discuss several approaches for parameterizing this win prediction model.
@@
2) Cost Threshold $\alpha\in[0,1]$: This threshold translates the predicted winning probability into a routing decision between $\mathcal{M}_{\text{weak}}$ and $\mathcal{M}_{\text{strong}}$. Given a query $q$, the routing decision is defined as:
$R^{\alpha}(q)=\begin{cases}\mathcal{M}_{\text{weak}}&\text{if }P(\text{win}_{s}\mid q)<\alpha,\\ \mathcal{M}_{\text{strong}}&\text{if }P(\text{win}_{s}\mid q)\geq\alpha.\end{cases}$ (2)
The threshold $\alpha$ controls the trade-off between quality and cost: a higher value of $\alpha$ enforces stricter cost constraints by favoring weak models more often, while a lower $\alpha$ biases toward higher-quality (but more expensive) strong models.
Finally, with the routing function $R^{\alpha}$ and two models, $M_{s}\in\mathcal{M}_{\text{strong}}$ and $M_{w}\in\mathcal{M}_{\text{weak}}$, we define a router model $M_{R^{\alpha}}:\mathcal{Q}\times\mathcal{M}_{\text{strong}}\times\mathcal{M}_{\text{weak}}\rightarrow\mathcal{A}$, which responds to a query $q$ as follows: [^2] For brevity, we denote this as $M_{R^{\alpha}}(q)$.
$M_{R^{\alpha}}(q,M_{s},M_{w})=\begin{cases}M_{s}(q)&\text{if }R^{\alpha}(q)=\mathcal{M}_{\text{strong}},\\ M_{w}(q)&\text{if }R^{\alpha}(q)=\mathcal{M}_{\text{weak}}.\end{cases}$ (3)
@@ 3.2 Metrics
In this section, we define evaluation metrics to capture the trade-off between cost and performance in the LLM routing problem. We begin with metrics that independently assess the cost efficiency and performance of a router model $M_{R^{\alpha}}$ routing between two models $M_{s}\in\mathcal{M}_{\text{strong}}$ and $M_{w}\in\mathcal{M}_{\text{weak}}$, and then introduce two compounded metrics used in our experimental evaluations.
@@
We measure the cost efficiency of $M_{R^{\alpha}}$ by calculating the percentage of calls to strong models:
$c(M_{R^{\alpha}})$ $=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathbb{I}\{R^{\alpha}(q)=\mathcal{M}_{\text{strong}}\},$ (4)
@@
since $\mathcal{M}_{\text{strong}}$ models are significantly more costly than $\mathcal{M}_{\text{weak}}$ models.
For performance, we calculate the average response quality:
$r(M_{R^{\alpha}})$ $=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}s(M_{R^{\alpha}}(q)),$ (5)
@@
where $s(M_{R^{\alpha}}(q))$ represents an individual response quality score, such as correctness on golden-labeled datasets (e.g., MMLU) or a numerical rating (e.g., 1-5 or 1-10), with higher values indicating better quality. Similarly, $r(M_{s})$ and $r(M_{w})$ can be defined for the strong and weak model respectively.
@@
Since the router model’s performance falls between that of the weak and strong models, we quantify its performance relative to the gap between them. We define the router’s overall performance improvement using the performance gap recovered (PGR):
$PGR(M_{R^{\alpha}})=\frac{r(M_{R^{\alpha}})-r(M_{w})}{r(M_{s})-r(M_{w})}.$ (6)
This metric captures how much of the performance difference between the weak and strong models is recovered by the router model.
@@
Neither of the above metrics alone is sufficient to capture the quality-cost trade-off in routing. For example, a trivial router that always sends queries to the strong model achieves a perfect $PGR=1$ but with no cost savings. To address this, we compute a call-performance graph for a router $M_{R^{\alpha}}$ by varying the threshold values $\alpha$. We then define the average performance gap recovered (APGR) as an overall measure of the router’s ability to recover the performance gap under different cost constraints:
$APGR(M_{R^{\alpha}})=\int_{0}^{1}PGR(M_{R^{\alpha}})\;d\left(c(M_{R^{\alpha}})\right).$ (7)
@@
In Figure 1-(right), APGR corresponds to the area between the router’s performance curve and the weak model’s performance curve. Empirically, we discretize the percentage of calls over the interval $[0\%,100\%]$ into $\{c_{i}\}_{i\in[10]}$. For each $c_{i}$, we determine the threshold $\alpha_{i}$ that satisfies the cost constraint. We approximate $APGR$ as:
$APGR(M_{R^{\alpha}})\approx\frac{1}{10}\sum_{i=1}^{10}PGR(M_{R^{\alpha_{i}}})$ (8)
@@
In many real-world applications, it is essential to quantify the cost required to achieve a certain level of performance. To address this, we introduce a second metric called call-performance threshold (CPT). Given a desired router performance, i.e., achieving a PGR of $x\%$, the CPT($x\%$) represents the minimum percentage of calls to the strong model needed to reach the desired PGR. In Figure 1-(right), the dotted green line illustrates CPT(50%), indicating the percentage of calls to GPT-4 needed to achieve a PGR of 50%. In this figure, $CPT(50\%)\approx 37\%$.